OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? 4 min 5 min 6 min

OpenStudy (lgbasallote):

use the equation \[\frac{t}{x} + \frac ty = 1\] where x and y are the time of the people let x = 10 minutes y = 15 minutes \[\frac{t}{10} + \frac{t}{15} = 1\] \[\frac{3t + 2t}{30} = 1\] \[\frac{5t}{30} = 1\] \[5t = 30\] solve for x

OpenStudy (lgbasallote):

i mean solve for t

OpenStudy (valpey):

Another way to say this is that X does one tenth of the job per minute and Y does one fifteenth of the job per minute. Working together they complete (one tenth plus one fifteenth equals) one sixth of the job per minute. When they work together for six minutes, they complete the job.